OpenStudy (anonymous):

If 4.88 grams of Zn reacts with 5.03 grams of S8 to produce 6.02 grams of ZnS, what are the theoretical yield and percent yield of this reaction? Be sure to show the work that you did to solve this problem. unbalanced equation: Zn + S8 yields ZnS Someone please help me!!!

OpenStudy (jfraser):

first and most important, balance the equation

OpenStudy (joannablackwelder):

Yep, @JFraser is right. Do you know how to balance, @firesquad ?

OpenStudy (anonymous):

Balanced equation: 8Zn + S8 ---> 8ZnS

OpenStudy (anonymous):

@JFraser @JoannaBlackwelder

OpenStudy (joannablackwelder):

Yep, good. Percent yield = (actual/theoretical)*100

OpenStudy (joannablackwelder):

Actual is the amount of product given in the problem.

OpenStudy (joannablackwelder):

Theoretical is the amount of product found using stoichiometry from the amount of reactant given.

OpenStudy (joannablackwelder):

*from the amount of limiting reactant

OpenStudy (anonymous):

I am stuck right here 4.88 g Zn x (1 mol Zn/ 65.38 g Zn) = 0.07464 mol Zn 5.03 g S8 x (1 mol S8/ 256.48 g S8) = 0.01961 mol S8

OpenStudy (joannablackwelder):

Ok, now use the molar ratios from the balanced reaction to convert to moles of ZnS

OpenStudy (anonymous):

what do i do after

OpenStudy (joannablackwelder):

For both calculations.

OpenStudy (anonymous):

would the ratio be 1: 8

OpenStudy (joannablackwelder):

For S8:ZnS, yes

OpenStudy (joannablackwelder):

But not for Zn:ZnS

OpenStudy (anonymous):

0.07464 mol Zn * (1/1)= 0.07464 mol Zn 0.01961 mol S8 * (8/1) = 0.1569 mol S8 is that how to do it

OpenStudy (joannablackwelder):

Well, that looks good for everything but the units at the end.

OpenStudy (anonymous):

What do I have to change for the units?

OpenStudy (joannablackwelder):

0.07464 mol ZnS 0.1569 mol ZnS

OpenStudy (joannablackwelder):

See why I did that?

OpenStudy (anonymous):

Oh Ok, now I find the percent yield right?

OpenStudy (joannablackwelder):

Well, we need to first determine which value of product to go with.

OpenStudy (joannablackwelder):

This is the limiting reactant part. Do you know how to determine that?

OpenStudy (anonymous):

The limiting reactant is S8

OpenStudy (joannablackwelder):

No, the limiting reactant is the reactant that produces less product.

OpenStudy (anonymous):

Sorry, it is Zn

OpenStudy (joannablackwelder):

Right :-)

OpenStudy (joannablackwelder):

So, what is the mass of product using the limiting reactant?

OpenStudy (anonymous):

Sorry for wait I had to do something. It is 0.07464 mol ZnS

OpenStudy (joannablackwelder):

No worries. That is the moles of product. What is the mass of product?

OpenStudy (anonymous):

65.38 g Zn

OpenStudy (joannablackwelder):

Hm, that's not what I get. How did you do that?

OpenStudy (anonymous):

I got it from the first formula

OpenStudy (joannablackwelder):

Hm, let me show you how I would do this step: 0.07464 mol ZnS(97.44 g/1 mol) = ? g ZnS

OpenStudy (anonymous):

Ok I get it

OpenStudy (joannablackwelder):

:-) So, can you find the percent yield?

OpenStudy (anonymous):

Wait 7.27 is the theoretical yield

OpenStudy (joannablackwelder):

Yep

OpenStudy (anonymous):

so what is the actual yield?

OpenStudy (joannablackwelder):

The amount of product given in the problem.

OpenStudy (anonymous):

Thank you so much

OpenStudy (joannablackwelder):

You're welcome! :-)